OpenStudy (anonymous):

The density of seawater is 1.03g/cm^3. what is the mass of 1 L of seawater?

OpenStudy (anonymous):

1L in cm^3 = ? is litre counted in m^3 ?

OpenStudy (anonymous):

Let \(\rho\) represent density, let \(m\) represent mass and let \(V\) represent volume \[\rho=\frac{1.03g}{cm^3}\times\frac{1000cm^3}{1L}=1030g/L\] We know \[\rho=\frac{m}{V}\] So therefore we can substitute: \[1030g/L=\frac{m}{1L}\rightarrow m=\frac{1030g}{1L}\times\frac{1L}{1}=1030g\] \(\therefore\) The mass of one litre of seawater is 1030 grams.

OpenStudy (anonymous):

THANK YOU SO MUCH :)

OpenStudy (anonymous):

Anytime Mary :)